OpenStudy (sgayton27):

Please help me solve the following : sqrt(7x+sqrt(7x+sqrt(7x)))

OpenStudy (confluxepic):

Welcome to Open Study.

OpenStudy (sgayton27):

\[\sqrt{7x+\sqrt{7x+\sqrt{7x}}}\]

OpenStudy (confluxepic):

Is there a value for x?

OpenStudy (sgayton27):

Thank you :)

OpenStudy (confluxepic):

Your'e welcome.

OpenStudy (confluxepic):

I am not sure if the equation is simplified or if the equation can be simplified.

OpenStudy (confluxepic):

It would be possible to solve it if there was a value for x.

OpenStudy (sgayton27):

wow I am so sorry I need to find the derivative

OpenStudy (xapproachesinfinity):

hehe solve what?

OpenStudy (xapproachesinfinity):

when you say solve you need something else to do the solving equality or inequality

OpenStudy (xapproachesinfinity):

Derivative! this unpleasant one hehe not that it's hard! first right thing down as rational powers

OpenStudy (xapproachesinfinity):

then do Chain Rule multiple times

OpenStudy (xapproachesinfinity):

your function is composed many times x==>7x===>root7x====>7x+root(7x)===>root(7x+root(7x))===>root of that again